TFY4104 Fysikk Eksamen onsdag 18. desember 2013 kl 0900 – 1300.

Kandidatnummer:

Oppgave A B C D Oppgave A B C D
1  26
2  27
3  28
4  29
5  30
6  31
7  32
8  33
9  34
10  35
11  36
12  37
13  38
14  39
15  40
16  41
17  42
18  43
19  44
20  45
21  46
22  47
23  48
24  49
25  50

NB: Kontroller at du har satt nøyaktig ETT kryss for HVER av de 50 oppgavene!